OpenStudy (jazzyluvsu123):

If a car can go from 0 to 60 mi/hr in 8.0 seconds, what would be its final speed after 5.0 seconds if its starting speed were 50 mi/hr?? PLEASE HELP ME WORK THIS OUT - i will fan you and give free medals

OpenStudy (anonymous):

Let's use Newton's Laws of Motion to determine a few things... First, the acceleration of the car: a=v f −v i t , where v f =26.82ms ;v i =0;t=8.0s a=26.82−08.0 =3.35ms 2 Now, to determine the final speed when v i =22.35ms ;t=5s;a=3.35ms 2 v f =v i +at v f =22.35+3.35⋅5=39.1ms The same problem in English units of miles per hour: Acceleration of the car from 0mi/h to 60mi/h in 8.0s: a=v f −v i t , where v f =60mi/h;v i =0;t=8.0s a=60mi/h−08.0s =7.5mi/h/s Final speed when initial velocity is 50mi/h: v i =50mi/h ; a=7.5mi/h/s ; t=5.0s v f =v i +at v f =50mi/h + (7.5mi/h/s )(5.0s ) v f =50mi/h+37.5mi/h=87.5mi/h